During a downhill ski race, the distance (in yds) a skier has travelled since leaving the starting gate follows this quadratic equation: D=5t^2+32t+1000 where t is the time in seconds measuring her run. After how many seconds has she skied 2,000yds?
Answer by stanbon(75887)

During a downhill ski race, the distance (in yds) a skier has travelled since leaving the starting gate follows this quadratic equation: D=5t^2+32t+1000 where t is the time in seconds measuring her run. After how many seconds has she skied 2,000yds?
-------------------------
5t^2+32t+1000 = 2000
5t^2 + 32t - 1000 = 0
-----
3t^2 + 16t - 500 = 0
-----
t = 11.30 seconds
-----------------
Cheers,
Stan H.
-----------------